A stone falls from a railroad overpass which is 36 ft high into the path of a train which is approaching the overpass with a uniform speed. If the stone falls when the train is 50 ft away from the overpass and the stone hits the ground just as the train arrives at that spot, how fast is the train moving?
Answer #1
find time to fall.

h=1/2 g t^2 find time. g=32ft/sec^2

Now, the train: d=rt

50ft=v*t

v will be in ft/sec
answered by: iyonna
Answer #2
d = Vo*t + 0.5gt^2 = 36 ft.
0 + 0.5*32t^2 = 36,
16t^2 = 36,
t^2 = 2.25,
t = 1.5s.

d = Vt,
V = d / t = 50 / 1.5 = 33.33 Ft / s.
answered by: yvteet